# appendix,by

### v1

@@ 4.3 $\mathsf{NTP}_{2}$ for existentially closed valued difference fields of residual characteristic zero
We now deviate from the proof of [CH14, Theorem 4.1]: instead of constructing $(a^{\omega}_{ij})_{i,j<\omega}$ and running their step (VI), we argue as follows. By Lemma˜4.38 and compactness there is a quantifier-free formula $\psi(x,c)\in\operatorname{qftp}(a/K)$ without symbols in $\mathrm{Mor}_{k}$ such that $\psi(\mathfrak{U},c)\subseteq\varphi(\mathfrak{U},a_{00})$. As $K=\bigcup_{\beta<\omega}a^{\beta}_{00}$ is an increasing union, we may assume $c=a_{00}^{\beta}$ for some $\beta<\omega$. Then $\psi(x,a^{\beta}_{00})$ is a quantifier-free formula, and by indiscernibility of the rows and strong homogeneity of $\mathfrak{U}$ we have $\psi(\mathfrak{U},a^{\beta}_{ij})\subseteq\varphi(\mathfrak{U},a_{ij}^{\beta})$ for all $i,j<\omega$, preserving inconsistency, while $a\vDash\psi(x,a_{i0}^{\beta})_{i<\omega}$ by indiscernibility of $(a^{\beta}_{i0})_{i<\omega}$ over $a$, so $(a^{\beta}_{ij})_{i,j<\omega}$ witnesses that $\psi$ has $\mathsf{TP}_{2}$. In particular, $\psi$ has $\mathsf{IP}$ witnessed by some $(a_{i})_{i<\omega}$ and $(b_{w})_{w\subseteq\omega}$. But then, writing $\psi(x,y)=\rho(x,\sigma(x),\ldots,\sigma^{n}(x),y,\sigma(y),\ldots,\sigma^{n}(y))$ for some $n$ and a formula $\rho$ in the language of $\mathcal{L}_{s,\iota}$-valued fields, we get that $\rho$, $(a^{\beta}_{i},\sigma(a^{\beta}_{i}),\ldots,\sigma^{n}(a^{\beta}_{i}))_{i<\omega}$ and $(b_{w},\sigma(b_{w}),\ldots,\sigma^{n}(b_{w}))_{w\subseteq\omega}$ witness that the $\mathcal{L}_{s,\iota}$–valued algebraically closed field $\mathfrak{U}$ has $\mathsf{IP}$. But algebraically closed $\mathcal{L}_{s,\iota}$-valued fields of residual characteristic zero are $\mathsf{NIP}$, as can be shown, e.g. by using [Kes24, Theorem 2.2]. This is a contradiction. ∎

### v2

@@ 4.3 $\mathsf{NTP}_{2}$ for existentially closed valued difference fields of residual characteristic zero
We now deviate from the proof of [CH, Theorem 4.1]: instead of constructing $(a^{\omega}_{ij})_{i,j<\omega}$ and running their step (VI), we argue as follows. By Lemma˜4.38 and compactness there is a quantifier-free $\mathcal{L}$-formula $\psi(x,c)\in\operatorname{qftp}(a/K)$ such that $\psi(\mathfrak{U},c)\subseteq\varphi(\mathfrak{U},a_{00})$. As $K=\bigcup_{\beta<\omega}a^{\beta}_{00}$ is an increasing union, we may assume $c=a_{00}^{\beta}$ for some $\beta<\omega$. Then $\psi(x,a^{\beta}_{00})$ is a quantifier-free formula, and by indiscernibility of the rows and strong homogeneity of $\mathfrak{U}$ we have $\psi(\mathfrak{U},a^{\beta}_{ij})\subseteq\varphi(\mathfrak{U},a_{ij}^{\beta})$ for all $i,j<\omega$, preserving inconsistency, while $a\vDash\psi(x,a_{i0}^{\beta})_{i<\omega}$ by indiscernibility of $(a^{\beta}_{i0})_{i<\omega}$ over $a$, so $(a^{\beta}_{ij})_{i,j<\omega}$ witnesses that $\psi$ has $\mathsf{TP}_{2}$. In particular, $\psi$ has $\mathsf{IP}$ witnessed by some $(a_{i})_{i<\omega}$ and $(b_{w})_{w\subseteq\omega}$. But then, writing $\psi(x,y)=\rho(x,\sigma(x),\ldots,\sigma^{n}(x),y,\sigma(y),\ldots,\sigma^{n}(y))$ for some $n$ and a formula $\rho$ in the language of $\mathcal{L}_{s,\iota}$-valued fields, we get that $\rho$, $(a^{\beta}_{i},\sigma(a^{\beta}_{i}),\ldots,\sigma^{n}(a^{\beta}_{i}))_{i<\omega}$ and $(b_{w},\sigma(b_{w}),\ldots,\sigma^{n}(b_{w}))_{w\subseteq\omega}$ witness that the $\mathcal{L}_{s,\iota}$–valued algebraically closed field $\mathfrak{U}$ has $\mathsf{IP}$. But algebraically closed $\mathcal{L}_{s,\iota}$-valued fields of residual characteristic zero are $\mathsf{NIP}$, as can be shown, e.g. by using [Kesting, Theorem 2.2]. This is a contradiction. ∎